OpenStudy (anonymous):

Create your own unique quadratic equation in the form y = ax^2 + bx + c. Use complete sentences and show all work to determine the following: Does the graph open up or down? How do you know? (2 points) Explain whether the graph has a maximum or minimum point. (2 points) Find the vertex and x-intercepts of the graph. (3 points)

OpenStudy (anonymous):

1.open up

OpenStudy (anonymous):

I have to make my own quadratic equation. I don't know how..

OpenStudy (anonymous):

well it determines on the quadratic equation if it is negative then down and if positive the up

sam (.sam.):

Simply make one, x^2-2x+1=0 open up because coefficient of x^2 is postivie

OpenStudy (anonymous):

Well, will you help me if I make one real quick? Let's answer this problem with this one. x^2 - 3x + 2 = 0.

OpenStudy (anonymous):

well i better start up with the y=x^2

OpenStudy (anonymous):

okay, y = x^2 - 3x + 2 = 0.

OpenStudy (anonymous):

quadratic equation - \[2x ^{2}+100x-3\]

OpenStudy (anonymous):

then to make it more complex add or subtract any number

sam (.sam.):

we'll take x^2 - 3x + 2 = 0 -open up because coefficient of x^2 is positive -Its minimum because coefficient of x^2 is positive -complete the square to find the vertex

OpenStudy (anonymous):

Okay we will use 2x^2 + 100x - 3.

OpenStudy (anonymous):

it has many maximum point but the minimum will about 3 or 4 or 5max

OpenStudy (anonymous):

because it is a parabola

OpenStudy (anonymous):

I don't know how to complete squares.

OpenStudy (anonymous):

Is there a formula?

sam (.sam.):

x^2 - 3x + 2 = 0 (x-3/2)^2-9/4+2=0 (x-3/2)^2-1/4=0 vertex at (3/2,1/4)

sam (.sam.):

vertex at (3/2,-1/4)

OpenStudy (anonymous):

What about the x intercept?

sam (.sam.):

factor x^2 - 3x + 2 = 0 (x-1)(x-2)=0 set each brackets =0 and solve for x

OpenStudy (anonymous):

So the x intercepts are (x-1) and (x-2)